If the sides of a square are lengthened by 4cm, the area becomes 256 square cm. Find the length of a side of the original square.
Thank you

Answer by stanbon(48569)

If the sides of a square are lengthened by 4cm, the area becomes 256 square cm. Find the length of a side of the original square.
Original square Data:
side=x; Area = x^2
New square Data:
side=x+4; Area = (x+4)^2
EQUATION:
New Area = 256 sq. cm.
(x+4)^2= 16^2
x+4= 16
x=12 cm. (length of side of original square)
Cheers,
Stan H.